Find value of sec and tan when sec+tan=4/3Also find value of sec*tan

Grade:9

1 Answers

Arun
6 years ago
Dear Adwait
 
hi ,
      Given , secx + tanx = 4/3,
           we know that 1+tan2 (x )= sec2 (x)
             =>     sec2 (x)-tan2 (x ) = 1
             =>     (sec (x) + tan (x))(sec (x) -tan(x)) = 1
                               sec (x) -tan(x) = ¾
                               sec (x) +tan(x) = 4/3
                        => sec (x) =25/24 & tan (X) = 7/24
                              sec (x) * tan (x) = 175/ 576
                                                                                     thank you !
 
Regards
Arun (askIITians forum expert)